# Multimodal Fusion Balancing Through Game-Theoretic Regularization

### v2

@@ Abstract
Multimodal learning can complete the picture of information extraction by uncovering key dependencies between data sources. However, current systems fail to fully leverage multiple modalities for optimal performance. This has been attributed to modality competition, where modalities strive for training resources, leaving some underoptimized. We show that current balancing methods struggle to train multimodal models that surpass even simple baselines, such as ensembles. This raises the question: how can we ensure that all modalities in multimodal training are sufficiently trained, and that learning from new modalities consistently improves performance? This paper proposes the Multimodal Competition Regularizer (MCR), a new loss component inspired by mutual information (MI) decomposition designed to prevent the adverse effects of competition in multimodal training. Our key contributions are: 1) Introducing game-theoretic principles in multimodal learning, where each modality acts as a player competing to maximize its influence on the final outcome, enabling automatic balancing of the MI terms. 2) Refining lower and upper bounds for each MI term to enhance the extraction of task-relevant unique and shared information across modalities. 3) Suggesting latent space permutations for conditional MI estimation, significantly improving computational efficiency. MCR outperforms all previously suggested training strategies and is the first to consistently improve multimodal learning beyond the ensemble baseline, clearly demonstrating that combining modalities leads to significant performance gains on both synthetic and large real-world datasets.
@@ 1 Introduction
Exploiting multimodal data has made significant progress, with advances in generalizable representations and larger datasets enabling solutions to previously unattainable tasks [27; 29; 37; 43; 42; 44; 50; 53; 62]. However, studies indicate that multimodal data is often utilized suboptimally, underperforming compared to ensemble unimodal models or even the best single modality [55; 60]. The expectation that adding a modality should improve performance, assuming independent errors and above-chance predictive power [17], is frequently contradicted in practice.
Huang et al. [20] attribute this issue to modality competition, where one modality quickly minimizes training error, misdirecting and suppressing the learning of others. Factors like noise levels, relationship complexity with the target, feature dimensionality, and data quality can cause one modality to fit faster than another. This imply that adding task-relevant information doesn’t guarantee better performance, primarily due to complications during training. To address these issues, it’s crucial to monitor each modality’s contribution during training and apply corrective measures.
Several balancing strategies have been proposed to tackle this issue [5; 6; 9; 10; 21; 26; 28; 40; 41; 54; 55; 57; 60]. A central aspect of these methods is estimating each modality’s contribution to the output. Most assume distributional independence between modalities on predicting the target, measuring contribution via unimodal performance [60; 26; 40; 6; 54]. Some methods bypass this assumption by estimating influence based on prediction differences between original and perturbed inputs [28; 21; 10]. Perturbations can take various forms, such as zeroing values [28], adding Gaussian noise [10], or using task-specific augmentations [21; 31]. These methods aim to amplify a modality’s influence by increasing the impact of perturbations on the output. However, this also makes the network more sensitive to these changes (e.g., noise), risks becoming overly reliant on the perturbations, and struggles to scale when multiple perturbations are required. Additionally, increasing the contribution of one modality can be achieved by overshadowing others, leading to an imbalance that undermines overall performance and makes the objective counterproductive.
@@
Given these challenges, how can we design an efficient regularization method that addresses multimodal competition, ensuring balanced and effective learning across all modalities?
Figure 1: (Left) Illustration of the conditional mutual information ($\operatorname{CMI}$) terms, $\operatorname{CMI}_{1}:I(X_{1};Y\mid X_{2})$ and $\operatorname{CMI}_{2}:I(X_{2};Y\mid X_{1})$, representing the unique contributions ($U_{1}$ and $U_{2}$) of each modality to the target. The shared task-relevant information ($S$) between the modalities is defined as $I(X_{1};X_{2})-I(X_{1};X_{2}\mid Y)$. (Right) Accuracy as a function of the ratio between the unique information ($U_{1}$) from modality $X_{1}$ and the shared information ($S$) between the modalities. Synthetic data are generated as $X_{1}=N_{1}+Y,X_{2}=N_{1}+Y$ where $N_{1},N_{2}$ are independent noise for each modality. We consider $S$ the percentage of the datapoints that both modalities have information about the label $Y$, $U_{1}$ and $U_{2}$ when only one has with the other modality equating to noise for those datapoints. In the experiment we keep $U_{2}$ constant while changing $U_{1}$ and $S$. As $U_{1}$ increases and $S$ decreases, accuracy deteriorates, reflecting intensified multimodal competition. Among the various methods, including Singleloss, Multiloss, Ensemble, unimodally pretrained and finetuned encoders (Uni-Pre Fine), OGM [40], AGM [28], and MLB [26]our regularization method MCR demonstrates a slower decline in accuracy. For further detals prease refer to Section 4.1
In this paper, we introduce the Multimodal Competition Regularizer ($\operatorname{MCR}$), a loss function designed to promote the exploration of task-relevant information across all available modalities. By decomposing the joint mutual information ($\operatorname{MI}$), we separately model shared and unique task-relevant information within the modalities. To efficiently capture the unique information from each modality, we employ a computationally inexpensive permutation-based approach. Our method maximizes the lower bounds of each MI term to encourage the network to learn both shared and unique information, while minimizing upper bounds on terms to suppress task-irrelevant information. We frame the problem in a game-theoretic setting, exploring strategies that involve both collaboration and competition among modalities to address the conflicting objectives that arise when increasing all modalities’ contributions simultaneously. This approach allows their contributions to adapt dynamically, achieving balance during training.
We extensively evaluate $\operatorname{MCR}$ on synthetic datasets and several established real-world multimodal benchmarks, including action recognition on AVE [49] and UCF [45], emotion recognition on CREMA-D [4], human sentiment on CMU-MOSI [61], human emotions on CMU-MOSEI [63] and egocentric action recognition on Something-Something [15]. Our results demonstrate that $\operatorname{MCR}$ is the first balancing method to significantly improve supervised multimodal training over the ensemble baseline across a variety of datasets and models. Our key contributions are summarized as follows:
@@
    An analysis of multimodal competition, defining the error increase caused in multimodal training, while demonstrating in our results that most previous methods do no outperform simple baselines, such as unimodal ensembles.
@@
    A novel multimodal training strategy, $\operatorname{MCR}$, designed to regularize multimodal competition, which includes:
    • Defining lower and upper bounds of the $\operatorname{MI}$ terms, encouraging the exploration of information across all modalities.
    • Introducing a game-theoretic perspective where modalities form the players that compete for training resources, assisting the regularization through balancing the corresponding MI terms.
    • Suggesting latent-space perturbations as an efficient way to estimate the lower bound of the $\operatorname{CMI}$ reducing the computational cost of multiple forward passes.
2 Analysis of Multimodal Competition
Consider a dataset of $N$ i.i.d. datapoints sampled from a distribution $\mathcal{D}$, where each datapoint consists of $M$ modalities $X=(X_{1},..,X_{M})$ and a target $Y$. Our objective is to learn a parameterized function $f:X;\theta\rightarrow Y$. We define the unimodal encoder for each modality as $f_{m}:X_{m};\theta_{m}\rightarrow Z_{m}$, which encodes input $X_{m}$ into a latent representation $Z_{m}$. The fusion network $f_{c}:[Z_{1},\dots,Z_{M}];\theta_{c}\rightarrow Y$ takes these latent representations and predicts the target $Y$, as does independently on each modality the unimodal task head $f_{c_{m}}:Z_{m};\theta_{c_{m}}\rightarrow Y_{m}$. The families of unimodal and multimodal models are defined as follows:
Unimodal models $\text{Unimodal models }\mathcal{F}_{u_{m}}:f_{u_{m}}\left(X_{m};\theta_{u_{m}}\right)=f_{c_{m}}\left(f_{m}\left(X_{m};\theta_{m}\right);\theta_{c_{m}}\right),\text{ for $m=[1,..,M]$.}$ (1)
$\text{Multimodal models }\mathcal{F}:\quad f(X;\theta)=f_{c}\left(f_{1}\left(X_{1};\theta_{1}\right),\dots,f_{M}\left(X_{M};\theta_{M}\right);\theta_{c}\right).$ (2)
@@ 2 Analysis of Multimodal Competition
The limitation of supervised multimodal training
In supervised learning with multimodal inputs $X_{1}$ and $X_{2}$, the objective is to learn representations $Z_{1}=f_{1}(X_{1};\theta_{1})$ and $Z_{2}=f_{2}(X_{2};\theta_{2})$ that, when combined through a fusion model $f_{c}(Z_{1},Z_{2})$, accurately predict a target $Y$. This can be expressed as minimizing the expected task loss or by maximizing the mutual information between the fused representation and the target:
$(Z_{1},Z_{2})=\mathop{\arg\max}_{\begin{subarray}{c}Z_{1}:=f_{1}(X_{1};\theta_{1}),\\ Z_{2}:=f_{2}(X_{2};\theta_{2})\end{subarray}}I(f_{c}(Z_{1},Z_{2});Y).$ (3)
During training, the model often undertrains the weaker modality, prioritizing the more accessible or higher-quality one. Such over-reliance on one modality limits the model’s ability to effectively utilize all available information. Consequently, the mutual information becomes dominated by the stronger modality, resulting in $I(f_{c}(Z_{1},Z_{2});Y)\approx I(Z_{1};Y)$. In this scenario, the conditional mutual information ($\operatorname{CMI}$) $I(Z_{2};Y\mid Z_{1})\approx 0$, indicating that once the model learns from $Z_{1}$, the information from $Z_{2}$ contributes little to predicting $Y$. Appendix A.1 presents a brief experiment demonstrating this phenomenon.
This phenomenon is not exclusive to multimodal learning. In single-modality feature learning, models often fit faster to dominant features, neglecting others that could enhance generalization. Regularization techniques like L1/L2 penalties and dropout [38; 46] were introduced to encourage balanced feature use and promote the learning of diverse patterns. While these techniques have become the standard in training models, their adaptation to multimodal learning has proven more challenging. For instance, Xiao et al. [58] attempted to apply dropout individually to each modality, but subsequent research [40] demonstrated that this approach is still limited in addressing the core issue. The central problem lies in effectively regulating how modalities interact and compete, which remains an open question in the field.
Multimodal competition
Multimodal competition occurs when the network primarily optimizes for one modality, leading to a decline in generalization. That modality reduces training error by overfitting to the data, limiting gradient feedback to the other modalities. Huang et al. [20] demonstrate that in late fusion models (i.e., $\theta_{c}=\emptyset$), each modality has a probability of causing multimodal competition. They define the phenomenon in terms of the correlation $\Gamma=\sup\langle X_{,}W\rangle$ of each input modalities and the weights of the output layer. If a modality dominate the competition it causes the other modality to maintain its initial (before training) correlation levels. We proceed this analysis by introducing the generalization error $\epsilon$ resulting from the effects of multimodal competition.
Definition of Multimodal Competition Error (MCE): Let $\Gamma_{1}$ and $\Gamma_{2}$ represent the correlations of modalities $X_{1}$ and $X_{2}$ with the output. If, at the end of training, $\Gamma_{1}\gg\Gamma_{2}$, and modality $X_{2}$ has predictive power greater than randomness ($R(X_{2})<R(\text{random})$), while making errors independent from those of $X_{1}$ ($\mathbb{E}[e_{1}e_{2}]=0$, where $e_{1}$ and $e_{2}$ are the errors of $X_{1}$ and $X_{2}$, respectively), then for a trained model $f$ and its optimal solution $f^{*}$, there exists a multimodal competition error $\epsilon$, such:
$\epsilon\geq R(f^{*})-R_{\text{emp}}(f^{*})\text{ and }\epsilon\leq R(f)-R_{\text{emp}}(f)\,$ (4)
where $R(f)=\mathbb{E}[L(y,f(x))]$ and $R_{\text{emp}}(f)$ denote the generalization and empirical model risks.
A higher $\epsilon$ indicates a stronger effect of multimodal competition, implying that the dominance of one modality significantly impacts the model’s generalization. In the following section, we introduce $\operatorname{MCR}$ as method to minimize $\epsilon$ by maximizing each modality’s correlation with the output.
@@ 3 Multimodal Competition Regularizer
Figure 2: Multimodal Competition Regularizer (MCR): The diagram illustrates the MCR framework used to mitigate competition between modalities in multimodal learning. Raw data ($X_{1}$ and $X_{2}$) are fed into the respective encoders, generating latent representations ($Z_{1}$ and $Z_{2}$). Each unimodal representation is permuted to create $\tilde{Z}_{1}$ and $\tilde{Z}_{2}$, and the three combinations are fed into the Fusion Network, producing predicted outputs ($Y$, $Y_{1}$, $Y_{2}$). The comparison between the predictions allows for an assessment of unimodal contribution. For instance, if $Y\approx Y_{1}$, then the disturbance of $Z_{1}$ had no impact on the result, indicating that the model relies solely on modality $X_{2}$, making the contribution of $X_{1}$ negligible. The MCR loss consists of three components: $\mathcal{L}_{{MIPD}1}$ and $\mathcal{L}{{MIPD}_{2}}$ maximize the Jensen-Shannon divergence (JSD) between the task output and predictions with permuted modalities, ensuring that each modality has the maximum information contribution. $\mathcal{L}_{\operatorname{Con}}$ encourages alignment between the representations of the modalities, promoting to capture the shared information. $\mathcal{L}_{\operatorname{CEB}}$ applies the conditional entropy bottleneck to regulate the mutual information between the combined representation $Z$ and the reconstructed output $\hat{Z}$. This framework aims to balance the contributions of each modality while minimizing competition, allowing for more robust multimodal representations.
In this section, we introduce $\mathcal{L}_{\operatorname{MCR}}$, a set of targeted loss components designed to directly address and mitigate the challenges of multimodal competition. These losses are derived by framing the multimodal learning task as optimizing the mutual information between the target $Y$ and the joint distribution of unimodal input features $X=(X_{1},X_{2})$. The method is focusing on two input modalities, though an extension to $M$ modalities is presented in Appendix A.5. The motivation behind these loss components stems from this formulation. To clearly define each loss, we first decompose the mutual information into distinct terms:

### v3

# Balancing Multimodal Training Through Game-Theoretic Regularization

@@ Abstract
Multimodal learning holds promise for richer information extraction by capturing dependencies across data sources. Yet, current training methods often underperform due to modality competition, a phenomenon where modalities contend for training resources leaving some underoptimized. This raises a pivotal question: how can we address training imbalances, ensure adequate optimization across all modalities, and achieve consistent performance improvements as we transition from unimodal to multimodal data? This paper proposes the Multimodal Competition Regularizer (MCR), inspired by a mutual information (MI) decomposition designed to prevent the adverse effects of competition in multimodal training. Our key contributions are: 1) A game-theoretic framework that adaptively balances modality contributions by encouraging each to maximize its informative role in the final prediction 2) Refining lower and upper bounds for each MI term to enhance the extraction of both task-relevant unique and shared information across modalities. 3) Proposing latent space permutations for conditional MI estimation, significantly improving computational efficiency. MCR outperforms all previously suggested training strategies and simple baseline, clearly demonstrating that training modalities jointly leads to important performance gains on both synthetic and large real-world datasets. We release our code and models at https://github.com/kkontras/MCR.
@@ 1 Introduction
Exploiting multimodal data has made significant progress, with advances in generalizable representations and larger datasets enabling solutions to previously unattainable tasks [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. However, studies indicate that jointly trained multimodal data is often utilized suboptimally, underperforming compared to ensembles of unimodal models, jointly trained modalities, or even the best single modality [11, 12]. The expectation that adding a new modality should improve performance, assuming independent errors and above-chance predictive power [13], is frequently contradicted in practice.
Figure 1: (Left) Illustration of the conditional mutual information ($\operatorname{CMI}$) terms, $\operatorname{CMI}_{1}:I(X_{1};Y\mid X_{2})$ and $\operatorname{CMI}_{2}:I(X_{2};Y\mid X_{1})$, representing the unique contributions ($U_{1}$, $U_{2}$) of each modality. The shared task-relevant information ($S$) is defined as $I(X_{1};X_{2})-I(X_{1};X_{2}\mid Y)$. (Right) Accuracy on a synthetic dataset designed to induce multimodal competition. We vary the ratio of unique information from modality 1 ($U_{1}$) to shared information ($S$), while keeping the contribution of modality 2 ($U_{2}$) constant. As the imbalance increases (moving right on the x-axis), the performance of most methods drops. The standard Joint Training (Singleloss) approach shows a steep decline, highlighting its vulnerability to modality competition where one modality dominates and suppresses the other. In contrast, our method, $\operatorname{MCR}$, demonstrates greater robustness by maintaining the highest accuracy and exhibiting the slowest performance degradation. See Section 4.1 for more details.
Huang et al. [14] attribute this issue to modality competition, where one modality quickly minimizes training error, misdirecting and suppressing the learning of others. To counteract this effect, monitoring each modality’s contribution during training and applying corrective measures is crucial. To this end, several balancing strategies have been proposed [15, 16, 17, 18, 19, 20, 21, 22, 23, 11, 24, 12, 25, 26, 27, 28, 29]. Some ignore a modality’s contribution beyond its independent unimodal performance, while others address this by measuring output differences under input perturbation, but at the cost of increased sensitivity to these perturbations and significant computational overhead. Moreover, it is crucial to examine whether enhancing one modality’s influence on the output does not come at the expense of others, as this could undermine overall performance.
@@
Given these challenges, how can we efficiently regularize multimodal competition to ensure balanced and effective learning across modalities?
This paper introduces a loss function encouraging the exploration of task-relevant information across modalities, the Multimodal Competition Regularizer ($\operatorname{MCR}$). The approach incorporates the following key contributions:
@@
    MI Bounds: We decompose joint mutual information into task-relevant shared and unique components, using refined lower and upper bounds to promote informative signals and suppress noise.
@@
    Game-Theoretic Modality Balancing: We frame modality interaction through a game-theoretic framework, allowing each modality to adjust its contribution throughout training.
  3. 3.
    Efficient CMI Estimation: We introduce latent-space perturbations for low-cost conditional MI estimation, avoiding repeated full-model passes.
We extensively evaluate $\operatorname{MCR}$ on synthetic datasets and several established real-world multimodal benchmarks, including action recognition on AVE [30] and UCF [31], emotion recognition on CREMA-D [32], human sentiment on CMU-MOSI [33], human emotions on CMU-MOSEI [34], and egocentric action recognition on Something-Something [35]. Our results demonstrate that $\operatorname{MCR}$ outperforms all previous methods and simple baselines across various datasets and models, improving multimodal supervised training.
2 Problem Analysis and Related Work
Consider a dataset of $N$ independent and identically distributed (i.i.d.) datapoints sampled from a distribution $\mathcal{D}$, where each datapoint has $M$ modalities $X=(X_{1},\dots,X_{M})$ and a target $Y_{t}$. Our goal is to learn a parameterized function $f:X;\theta\rightarrow Y_{t}$, where $\theta$ denotes all the model’s learnable parameters. The unimodal encoder for each modality is defined as $f_{m}:X_{m};\theta_{m}\rightarrow Z_{m}$, encoding input $X_{m}$ into a latent representation $Z_{m}$. The fusion network $f_{c}:[Z_{1},\dots,Z_{M}];\theta_{c}\rightarrow Y_{t}$ predicts $Y_{t}$ from the latent representations, as do the unimodal task heads $f_{c_{m}}:Z_{m};\theta_{c_{m}}\rightarrow Y_{m}$. Model families are defined as, unimodal models for $m=[1,..,M]$ modalities:
$\mathcal{F}_{u_{m}}:f_{u_{m}}\left(X_{m};\theta_{m},\theta_{c_{m}}\right)=f_{c_{m}}\left(f_{m}\left(X_{m};\theta_{m}\right);\theta_{c_{m}}\right),$ (1)
and for multimodal models:
$\mathcal{F}:f(X;\theta)=f_{c}\left(\left[f_{1}\left(X_{1};\theta_{1}\right),..,f_{M}\left(X_{M};\theta_{M}\right)\right];\theta_{c}\right).$ (2)
@@ 2 Problem Analysis and Related Work
2.1 The limitation of supervised multimodal training
In supervised learning, the goal is to learn representations $Z_{1}=f_{1}(X_{1};\theta_{1})$ and $Z_{2}=f_{2}(X_{2};\theta_{2})$ when fused via $f_{c}(\left[Z_{1},Z_{2}\right])$, yield accurate predictions. This is achieved by minimizing the task loss or, equivalently, by maximizing the MI between the fused representation and the target:
$\mathop{\arg\max}_{\begin{subarray}{c}Z_{1}:=f_{1}(X_{1};\theta_{1}),\\ Z_{2}:=f_{2}(X_{2};\theta_{2})\end{subarray}}I(f_{c}(\left[Z_{1},Z_{2}\right]);Y_{t}).$ (3)
During training, models often over-rely on the stronger or more accessible modality, limiting the contribution of others. This leads to mutual information being dominated by one modality, e.g., $I(f_{c}(\left[Z_{1},Z_{2}\right]);Y)\approx I(Z_{1};Y)$ with $I(Z_{2};Y\mid Z_{1})\approx 0$, indicating that $Z_{2}$ adds little once $Z_{1}$ is learned. See Appendix A.1 for an illustrative experiment and Appendix A.2 for a formal definition of the resulting generalization gap.
This kind of imbalance is well-known in single-modality learning, where dominant features can overshadow others, harming generalization. Regularization techniques like $l_{1}\,/\,l_{2}$ penalties and dropout promote balanced feature use [36, 37], but their adaptation to multimodal settings is nontrivial. For example, applying modality-specific dropout [38] offers limited benefits [22]. The core challenge remains: how to effectively regulate interaction and competition between modalities.
2.2 Related Work
Prior research has explored various strategies for multimodal learning, ranging from simple unimodal and ensemble-based approaches to more sophisticated methods for balancing modality contributions. Unimodal training optimizes each modality separately, while ensemble methods combine unimodal predictions without additional training. Joint training optimizes all modalities under a single-loss objective but does not explicitly ensure sufficient training for each modality. To address this, Multi-Loss [39] introduces additional unimodal task losses, and MMCosine [40] equalizes modality influence by standardizing features and weights. Pre-trained unimodal encoders are often used, either with frozen weights (Uni-Pre Frozen) or fine-tuned jointly (Uni-Pre Finetuned). Other adaptive strategies include MSLR [12], which adjusts learning rates based on unimodal validation performance, OGM [22], which modulates gradients by comparing unimodal performance across modalities, and MLB [20], which combines unimodal task losses and modulates gradients from both unimodal and multimodal objectives, MMPareto [25] that mitigates gradient conflicts between modalities by equalizing the contribution of unimodal and multimodal gradient and D&R [26] suggest a new strategy where modalities that overfit get their part of the network partially reweighted with the initial weights of the training.
Most of these methods assume distributional independence and measure modality contributions through unimodal performance, which can be a limited indicator, missing cases where modality correlation is crucial. Other approaches estimate influence based on prediction differences after perturbations [21, 19, 18]. AGM [21] uses zero-masking Shapley values directly optimizing them as unimodal predictors, Wei et al. [29] use a permutation-based Shapley values and resampling of the training set to affect the training, while other methods address similar problems by introducing perturbations such as Gaussian noise [18] or task-specific augmentations [19, 41]. However, perturbation-based approaches increase the network’s sensitivity to the chosen perturbations and hinder scalability due to their higher computational demands.
A line of work keeps unimodal training as the primary strategy. MLA [27] uses a shared task head and dynamic weighted summation during validation, while ReconBoost [28] alternates unimodal updates with agreement and diversity regularization before finetuning the ensemble. However, these approaches avoid multimodal training in the earlier steps to mitigate conflicts, yet overlook the potential benefits of direct multimodal interactions in those steps.
@@ 3 Multimodal Competition Regularizer
Figure 2: Multimodal Competition Regularizer ($\operatorname{MCR}$): The diagram illustrates the $\operatorname{MCR}$ framework, which mitigates modality competition in multimodal learning. Raw data ($X_{1}$ and $X_{2}$) are encoded into latent representations ($Z_{1}$ and $Z_{2}$), which are then permuted to create $\tilde{Z}_{1}$ and $\tilde{Z}_{2}$ and the paired combinations. These combinations are passed through the Fusion Network to produce predicted outputs ($Y$, $\tilde{Y}_{1}$, $\tilde{Y}_{2}$). The comparison between predictions reveals each modality’s contribution. For example, if $Y\approx\tilde{Y}_{1}$, it shows that $X_{1}$ has little impact, and the model relies on $X_{2}$. The $\operatorname{MCR}$ loss includes three components: $\mathcal{L}_{\operatorname{MIPD}}$ maximize the Jensen-Shannon divergence (JSD) between task output and permuted modality predictions. $\mathcal{L}_{\operatorname{Con}}$ aligns modality representations, while $\mathcal{L}_{\operatorname{CEB}}$ penalizes task-irrelevant information by reconstructing back to the latent space.
Multimodal competition arises when a model trained on multiple modalities prioritizes one, leading to over-reliance and reducing the contribution of others. This imbalance limits the model’s ability to fully utilize all available information. In this section, we introduce $\mathcal{L}_{\operatorname{MCR}}$, a set of loss components designed to address multimodal competition. Each component of the loss is motivated by the following MI decomposition: